# references.bib

### v2

@@ 6 Acknowledgments
Appendix A Mathematical comparison of PPI and the method of Olofsson
We show that when the number of map product points $N$ is large, the PPI and Olofsson area estimation confidence intervals are similar. In particular, we show (1) that they have approximately the same length when the number of samples is large and (2) they are centered on the same point. These two facts together provide a mathematical explanation for why PPI and the Olofsson method approximately agree in all of our experiments.
We begin with notation. Estimating the area proportion of a land cover class in a given region is equivalent to estimating the mean of a binary variable $Y$, where $Y=1$ for points in the class and $Y=0$ for points outside the class. Suppose we have $n$ uniformly random sampled ground truth points $Y_{1},Y_{2},\dots Y_{n}$ and $N$ map product points $\hat{Y}_{1},\hat{Y}_{2},\dots\hat{Y}_{N}$. For all $0\leq i,j\leq 1$, let $n_{ij}$ denote the number of ground truth points that are in both map class $i$ and ground truth class $j$. Let $n_{i\cdot}=n_{i0}+n_{i1}$ denote the total number of ground truth points in map class $i$, and $n_{\cdot i}=n_{0i}+n_{1i}$ denote the total number of ground truth points in ground truth class $i$. Furthermore, let $A_{i}$ denote the fraction of the map product that is in class $i$. We have $A_{0}+A_{1}=1$. The map-imputed area estimate, which we will correct for bias, is $A_{1}$. We will use the approximations $A_{0}\approx\frac{n_{0\cdot}}{n}$ and $A_{1}\approx\frac{n_{1\cdot}}{n}$. This is a reasonable approximation since the $n$ ground truth locations are sampled uniformly at random.
@@ Appendix A Mathematical comparison of PPI and the method of Olofsson
We compare this with the Olofsson standard error
$SE_{\text{Olofsson}}$ $=\sqrt{A_{0}^{2}\frac{\frac{n_{00}}{n_{0\cdot}}\frac{n_{01}}{n_{0\cdot}}}{n_{0\cdot}}+A_{1}^{2}\frac{\frac{n_{10}}{n_{1\cdot}}\frac{n_{11}}{n_{1\cdot}}}{n_{1\cdot}}}$
@@
The expressions we arrive at for PPI and the method of Olofsson agree, indicating that for large sample sizes the standard errors are approximately equal. This implies that the confidence intervals from the two methods will have the same width.
@@
Next, we show that the PPI and Olofsson confidence intervals are centered around the same point. Turning first to the PPI point estimate
@@
The Olofsson area point estimate is
$\hat{\theta}_{\text{Olofsson}}$ $=A_{0}\frac{n_{01}}{n_{0\cdot}}+A_{1}\frac{n_{11}}{n_{1\cdot}}$

### v3

@@ 6 Acknowledgments
Appendix A Mathematical comparison of PPI and post-stratified area estimator
We show that when the number of map product points $N$ is large, the PPI and post-stratified area estimator confidence intervals are similar. In particular, we show (1) that they have approximately the same length when the number of samples is large and (2) they are centered on the same point. These two facts together provide a mathematical explanation for why PPI and the post-stratified estimator approximately agree in all of our experiments.
We begin with notation. Estimating the area proportion of a land cover class in a given region is equivalent to estimating the mean of a binary variable $Y$, where $Y=1$ for points in the class and $Y=0$ for points outside the class. Suppose we have $n$ uniformly random sampled ground truth points $Y_{1},Y_{2},\dots Y_{n}$ and $N$ map product points $\hat{Y}_{1},\hat{Y}_{2},\dots\hat{Y}_{N}$. For all $0\leq i,j\leq 1$, let $n_{ij}$ denote the number of ground truth points that are in both map class $i$ and ground truth class $j$. Let $n_{i\cdot}=n_{i0}+n_{i1}$ denote the total number of ground truth points in map class $i$, and $n_{\cdot i}=n_{0i}+n_{1i}$ denote the total number of ground truth points in ground truth class $i$. Furthermore, let $A_{i}$ denote the fraction of the map product that is in class $i$. We have $A_{0}+A_{1}=1$. The map-only area estimate, which we will correct for bias, is $A_{1}$. We will use the approximations $A_{0}\approx\frac{n_{0\cdot}}{n}$ and $A_{1}\approx\frac{n_{1\cdot}}{n}$. This is a reasonable approximation since the $n$ ground truth locations are sampled uniformly at random.
@@ Appendix A Mathematical comparison of PPI and post-stratified area estimator
We compare this with the post-stratified estimator standard error
$SE_{\text{post}}$ $=\sqrt{A_{0}^{2}\frac{\frac{n_{00}}{n_{0\cdot}}\frac{n_{01}}{n_{0\cdot}}}{n_{0\cdot}}+A_{1}^{2}\frac{\frac{n_{10}}{n_{1\cdot}}\frac{n_{11}}{n_{1\cdot}}}{n_{1\cdot}}}$
@@
The expressions we arrive at for PPI and the post-stratified estimator agree, indicating that for large sample sizes the standard errors are approximately equal. This implies that the confidence intervals from the two methods will have the same width.
@@
Next, we show that the PPI and post-stratified confidence intervals are centered around the same point. Turning first to the PPI point estimate
@@
The post-stratified area point estimate is
$\hat{\theta}_{\text{post}}$ $=A_{0}\frac{n_{01}}{n_{0\cdot}}+A_{1}\frac{n_{11}}{n_{1\cdot}}$